OpenStudy (anonymous):

Evaluate the function f(x)=3x-2,when x=(-1) A.-9 B.-5 C.1 D.5 Please help me to find answer and explain the steps to getting the answer.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do \[3\times (-1)-2\]

OpenStudy (anonymous):

-5 would be my answer ,correct?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

3 x -1 = -3 -3 - 2 is the same as -3 + -2 = -5

OpenStudy (misty1212):

yes it would \[\color\magenta\heartsuit\]